After the recent Aqua BB WUCD on 01/31/2015 it was observed that 3 detectors (2, 3,& 9) for band 21 experienced a b1 jump and NEdT increase. The increased NEdT is greater than the specification and has been consistent for over two weeks. MCST procedure dictates a QA LUT update for these detectors to ‘Noisy’ status. This update will be included in the next regular Aqua LUT delivery. The timestamp for the QA change will be 2015031.2100.
SFE Anomaly: After passage through the SAA on 2014/299 ~01:35 EDOS observed corrupted MODIS data. The cause was traced to a formatter card that had become out of sync resulting in the corrupt data. Only MODIS data is processed by this card so no other instruments were affected. FOT successfully reset the card and returned data flow to nominal operation. Expect a data loss from 2014/299 01:35:00 – 18:47:00. Exact loss time still to be confirmed by FOT.
Terra MODIS Collection 6 Level 1B products will be reprocessed to correct for the long term drift observed in the reflectance trending of band 5. More details regarding the reprocessing schedule can be found here: http://ladsweb.nascom.nasa.gov/news/news080414.html
Also, information regarding this calibration change can be found here.
The presentations from the MODIS Science Team Meeting, April-2014 are available at the following location: http://modis.gsfc.nasa.gov/sci_team/meetings/201404/
Terra experienced an anomaly for B36/D7 beginning on 2014109.1405. Unlike previous instances of this unknown electronics/formatter anomaly, the detector behavior has not returned to normal.After 2 weeks continually being in the anomaly state, the decision was made to change the QA flag for this detector from 'Noisy' to 'Inoperable'. All Band 36 detectors are currently flagged as Noisy.
This QA change will be in effect from the LUT versions 184.108.40.206 (Collection 5) and 220.127.116.11 (Collection 6)